OpenStudy (anonymous):

Find the arc length (L) of \[y^2=4(x+3)^3\] for the closed interval [0,1] (that'll mean a=0 & b=1). Formula for arc length. \[ L = \int\limits_{a}^{b} \sqrt{1+(\frac{dy}{dx})^2} dx \] The problem is I get either a messy cube root fraction in terms of Y, or a \(\pm\)root mess (kind of like what you get for circles, ellipses, or hyperbolas). Can somebody show me the steps for the differentiation into integration here? I assume it's probably some tip or trick I've over looked. (+1 medal will be given)

OpenStudy (anonymous):

hi agentx5 i dont know if this is right or not can we write \[y=\pm 2(x+3)^{3/2}\\ \frac{dy}{dx}=\pm 3(x+3)^{1/2} \\ (\frac{dy}{dx})^2= 9(x+3)\\ then\\ L=2 \int\limits_{0}^{1} \sqrt{1+9(x+3)} dx\] ?

OpenStudy (anonymous):

@agentx5

OpenStudy (anonymous):

\[2\cdot [\frac{9x^2}{2}+28x+C]_0^1\] \[2\cdot ((\frac{9(1)^2}{2}+28(1))-(\frac{9(0)^2}{2}+28(0))\] Hmm the correct answer should be \(\large\frac{104}{3}\)

OpenStudy (anonymous):

Nobody? :-/

OpenStudy (anonymous):

i will totally work on it after my exam :)

OpenStudy (anonymous):

Gracias/Thank You/Schönen Dank :-)

OpenStudy (anonymous):

\[L=\int_{0}^{1}\sqrt{1+\left[f'(x)\right]^2}dx\] \[y^2=4(x+3)^3=f(x)^2 \Rightarrow f(x)=\pm \sqrt{4(x+3)^3}=\pm 2\sqrt{(x+3)^3}\] \[f'(x)=\pm 3\sqrt{x+3}\] \[L=\int_{0}^{1}\sqrt{1+9(x+3)}dx=\int_0^1\sqrt{9x+4}dx=\frac{2}{27}(13\sqrt{13}-8)\] I'm not getting \(\frac{104}{3}. . .\)

OpenStudy (turingtest):

I see nothing wrong with @mukushla 's or @Limitless 's answers, so there is a typo somehwere

OpenStudy (turingtest):

oh wait, Limitless messed up

OpenStudy (turingtest):

\[L=\int_0^1\sqrt{1+9(x+3)}dx=\int_0^1\sqrt{9x+28}dx\]

OpenStudy (turingtest):

still don't get 104/3

OpenStudy (anonymous):

@mukushla did, as well.

OpenStudy (anonymous):

I don't see any reason for the \(2\) outside of the integration.

OpenStudy (turingtest):

oh yeah, neither do I

OpenStudy (turingtest):

but I think we all agree that the answer to the question as it is given is not 104/3

OpenStudy (anonymous):

Indeed, @TuringTest.

OpenStudy (anonymous):

@TuringTest y=2(x+3)^(3/2) and y=-2(x+3)^(3/2) thats why i put 2 outside is that right or ...

OpenStudy (anonymous):

@mukushla, the two disappears after simplification of the derivative.